### v1

@@ I Introduction
Recent studies have explored a diverse range of computational and intelligent approaches to improve the accuracy and efficiency of stock market forecasting. These works span from deep learning architectures and hybrid statistical-machine learning frameworks to quantum-inspired algorithms.
Aşırım, et al.[11171624] employed a single-layer Transformer model to forecast daily closing prices of Nasdaq stocks. By segmenting and independently normalizing historical data with sinusoidal positional embeddings, the model achieved an average correlation of about 0.96, effectively and efficiently capturing stock price trends.
Ahire et al.[10882429] proposed a hybrid Auto Regressive Integrated Moving Average (ARIMA) and machine learning model to forecast stock trends using ICICI and SBI data. By combining ARIMA residuals with technical indicators such as Relative Strength Index (RSI), Simple Moving Average (SMA), and Stochastic K, their method improved prediction accuracy to about 0.96 for SBI and 0.94 for ICICI, effectively identifying bullish and bearish movements.
Singh et al.[10545132] applied multiple machine learning algorithms to predict stock performance of HRM firms Info Edge Ltd and Quess Corporation using five years of data. Results showed that Random Forest and Gradient Boosting achieved the highest accuracy, with $R^{2}$ values above 0.999 and Mean Absolute Percentage Error (MAPE) below 0.2%.
RaviTeja et al.[10690751] developed a Novel Quantum Enforcing Algorithm (NQEA) to forecast stock index and Bitcoin prices, comparing it with the Bernstein-Vazirani Algorithm (BV). Using datasets with 1,400 training and 1,856 testing samples. Results showed that NQEA achieved an accuracy of 95.9% with a loss of 4.1%, outperforming BV’s 88.0% accuracy and 12.0% loss.
Hossain et al.[11171862] applied Linear Regression, Random Forest, Gradient Boosting, XGBoost, and Multilayer Perceptron (MLP), to predict stock prices of Dhaka Stock Exchange. Results showed that Linear Regression achieved the lowest prediction error and most consistent performance across all datasets, outperforming ensemble and neural network models.
@@
The rest of this article is organized as follows. Section II reviews the task and dataset used in this work. Section III details the specific architecture of our proposed QTCNN model. Section IV describes the experimental setup and presents our numerical results. Finally, Section V concludes this paper and discusses potential directions for future research.
@@ II-A JPX Tokyo Stock Exchange Prediction
II-B Task Overview
  • Prediction task. For each trading day, the model produces a cross-sectional ranking of stocks. Based on this ranking, we form a long–short portfolio by going long the top $N$ stocks and short the bottom $N$ stocks.
  • Model output. On each trading day, the model outputs a continuous score (or probability) for every stock. These scores are standardized cross-sectionally (within the day) and then used to rank the stocks.
  • Performance metric. We evaluate the model by computing the Sharpe ratio of the long–short portfolio over an out-of-sample (OOS) period, using the daily portfolio returns implied by the official competition target.
II-C Original Dataset and Stock Universe
II-C1 Data source and time span
@@
II-C2 Universe definition and Universe0
@@
II-D Key Fields and Financial Interpretation
II-D1 Price data (stock_prices)
The main price table (both in the training and supplemental sets) contains the following fields used in this work:
  • Date: Trading date.
  • SecuritiesCode: Stock identifier (e.g., 1301, 1332), serving as a key together with Date.
  • Open, High, Low, Close: Daily open, high, low, and close prices.
  • Volume: Daily traded volume (in shares).
  • AdjustmentFactor: Adjustment factor capturing stock splits, reverse splits, share distributions, and other corporate actions.
  • SupervisionFlag: Indicator for stocks under special supervision by the exchange or at risk of delisting.
  • Target: The official competition target, corresponding to a two-day holding return defined on future prices (see Section II-E).
@@ II-D1 Price data (stock_prices)
II-D2 Static stock information (stock_list)
The stock_list table provides relatively static attributes for each stock:
  • SecuritiesCode: Stock identifier, matching stock_prices.
  • Name: Stock name.
  • MarketCapitalization: Market capitalization.
  • Universe0: Boolean flag indicating membership in the core prediction universe.
@@ II-D2 Static stock information (stock_list)
II-D3 Adjusted close price
@@
II-E Definition and Interpretation of the Target
@@
II-F Derived Features and Preprocessing
II-F1 Single-stock time-series features
@@
II-F2 Cross-sectional normalization
@@
II-G Label Construction and Weakly Supervised Training Samples
@@
II-H Temporal Split and Sequence Construction
II-H1 Temporal train–test split
@@
II-H2 Sequence construction for time-series models
@@
II-I Subset Sampling of Trading Days
@@ II-I1 Stride sampling
II-I2 Year-wise stratified fraction sampling
@@
II-I3 Relation between subsets and the full dataset
@@ III-E Quantum Temporal Convolutional Neural Network (QTCNN)
The central methodological contribution of this work is the Quantum Temporal Convolutional Neural Network (QTCNN), which addresses two key limitations of existing quantum approaches for sequential data:
@@
    Parameter efficiency: Through intra-layer parameter sharing, QTCNN reduces quantum circuit parameters from $O(n_{q}\cdot L)$ to $O(L)$, mitigating trainability issues associated with barren plateaus in deep variational circuits.
@@ III-F5 Training Objective
III-G Quantum Long Short-Term Memory (QLSTM)
@@ III-G1 Sequence Embedding
III-G2 Quantum Gate Computation
@@
III-G3 LSTM Update Equations
@@
III-H Variational Quantum Neural Network (QNN)
@@ III-H3 Output Layer
III-I Quantum Kernel Support Vector Machine (QSVM)
@@ III-I3 SVM Optimization
III-J Quantum-Adaptive Self-Attention (QASA)
@@ III-J1 Transformer Encoder
III-J2 Quantum Enhancement Layer
@@ III-J3 Classification Output
III-K PQC-Enhanced Gradient Boosting (PQC+LGBM)
@@ III-K1 Stage 1: PQC Embedding
III-K2 Stage 2: LambdaRank Optimization
@@
III-L Classical Baseline Models
III-L1 Multi-Layer Perceptron (MLP)
@@
III-L2 Long Short-Term Memory (LSTM)
@@ III-L4 LightGBM Ranker
III-L5 Momentum-Volatility Baseline
@@
III-M Training Configuration
III-M1 Optimization Details
@@
III-M2 Hyperparameter Summary
TABLE I: Training Configuration Summary
@@ III-M2 Hyperparameter Summary
III-N Evaluation Framework
III-N1 Portfolio Construction
@@ III-N3 Statistical Significance
III-O Quantum Implementation Details
III-O1 Circuit Simulation
@@
III-O2 Hardware-Efficient Ansatz
@@ III-O3 Gradient Computation
III-P Reproducibility
III-P1 Random Seeds
@@ III-Q Conclusion
III-R Computational Considerations for Quantum Simulation
A critical challenge in variational quantum algorithm research is the exponential scaling of classical simulation. For an $n_{q}$-qubit system, state vector simulation requires $O(2^{n_{q}})$ memory and $O(L\cdot 4^{n_{q}})$ time complexity per forward pass, where $L$ denotes circuit depth. Table II reports the empirical wall-clock time for a single forward-backward pass across all quantum architectures evaluated in this study, measured on Pannylane.
TABLE II: Computational Cost per Training Iteration (on RTX3090 GPU and 12th Gen Intel(R) Core(TM) i9-12900KS CPU)
@@ IV-B Out-of-Sample Performance Analysis
Table III presents the out-of-sample (OOS) Sharpe ratio for all evaluated models over a 22-day testing period. Among all approaches, the quantum models demonstrate a clear advantage in terms of risk-adjusted returns.
@@
Classical/Quantum Model Sharpe Ratio (OOS) OOS Days
@@
Quantum QNN 0.467 $\pm$ 0.038 22
Quantum QCNN 0.361 $\pm$ 0.298 22
Quantum QLSTM 0.333 $\pm$ 0.051 22
Classical Transformer 0.313 $\pm$ 0.036 22
Quantum QASA 0.266 $\pm$ 0.045 22
Quantum QSVM 0.131 $\pm$ 0.129 22
Quantum PQC+LGBM 0.122 $\pm$ 0.017 22
Classical LSTM 0.044 $\pm$ 0.029 22
TABLE III: Out-of-sample performance summary aggregated from project outputs.
@@ IV-C Ablation Analysis
The performance hierarchy among quantum models provides insight into architectural design principles:

### v2

@@ I Introduction
TABLE I: Comparison of Recent Stock Prediction Methods
Study Method / Model Target Data Key Techniques Performance
Aşırım, et al.[11171624] Single-layer Transformer Nasdaq daily closing prices Data segmentation, independent normalization, sinusoidal positional embeddings Average correlation $\approx 0.96$, effectively captured stock price trends
Ahire et al.[10882429] Hybrid ARIMA + Machine Learning ICICI and SBI stock data ARIMA with RSI, SMA, and Stochastic K indicators Prediction accuracy $\approx 0.96$ (SBI) and $0.94$ (ICICI); effective bullish/bearish detection
Singh et al.[10545132] Random Forest, Gradient Boosting TCS, HCL, and Ques Corp (5-year data) Multi-model performance comparison $R^{2}>0.99$ and MAPE $0.2\%$, highest accuracy
RaviTeja et al.[10690751] Novel Quantum Enforcing Algorithm (NQEA) Stock index and Bitcoin prices Quantum-inspired optimization, compared with BV algorithm Accuracy $95.9\%$, loss $4.1\%$, outperforming BV (88.0%)
Hossain et al.[11171862] Linear Regression, RF, GB, XGBoost, MLP Dhaka Stock Exchange Cross-model performance comparison Linear Regression achieved lowest prediction error and most consistent performance
@@
Recent studies (shown as TABLE I) have explored a diverse range of computational and intelligent approaches to improve the accuracy and efficiency of stock market forecasting. These works span from deep learning architectures and hybrid statistical-machine learning frameworks to quantum-inspired algorithms.
@@
The rest of this article is organized as follows. Section II reviews the task and dataset used in this work. Section III details the architecture and provides a quantum resource complexity analysis of our proposed QTCNN model. Section IV presents numerical results including a qubit scalability study, mechanism analysis of quantum enhancement, and noise resilience experiments. Finally, Section V discusses dequantization considerations, prospects for NISQ and fault-tolerant execution, and future research directions.
TABLE II: Fields in the Price Data Table (stock_prices)
Field Name Description
Date Trading date.
SecuritiesCode Unique stock identifier (e.g., 1301, 1332), used as a key together with Date.
Open Daily opening price.
High Daily highest price.
Low Daily lowest price.
Close Daily closing price.
Volume Daily traded volume (in shares).
AdjustmentFactor Cumulative adjustment factor accounting for stock splits, reverse splits, dividends, and other corporate actions.
SupervisionFlag Indicator for stocks under exchange supervision or at risk of delisting.
Target Official competition target; two-day forward return based on future prices (see Section II-E).
TABLE III: Fields in the Static Stock Information Table (stock_list)
Field Name Description
SecuritiesCode Stock identifier matching stock_prices.
Name Official stock name.
MarketCapitalization Market capitalization of the company.
Universe0 Boolean flag indicating whether the stock belongs to the core prediction universe.
Figure 1: Overview of the Stock Prediction Task and Portfolio Evaluation Pipeline.
@@ II Task and Dataset
II-A Task Overview
FIGURE 1 illustrates the overall pipeline of the stock prediction task and portfolio evaluation framework used in this study. The process consists of data preparation, feature and label construction, model prediction, cross-sectional ranking, and portfolio performance evaluation.
Data and features. We begin with the JPX stock dataset, which provides daily price and volume information for all listed securities within the official prediction universe. From the raw price series, we construct a set of technical and statistical features, including returns, momentum indicators, volatility measures, and price–volume related variables. To ensure comparability across stocks on the same trading day, all features are standardized in a cross-sectional manner (within each day) using z-score normalization.
Label construction. For each stock $k$ on trading day $t$, the supervision signal is derived from the official competition target, which corresponds to the forward return over the next holding period. These targets serve as the ground-truth signals that guide the model to learn cross-sectional differences in expected future performance.
Model prediction. The engineered features for each day are fed into the predictive model (e.g., QTCNN or other baseline models), which outputs a continuous score $\hat{y}_{t}^{(k)}$ for every stock. This score reflects the model’s assessment of the relative attractiveness of each stock for the upcoming holding period.
Cross-sectional ranking and portfolio formation. On each trading day, stocks are sorted based on the predicted scores $\hat{y}_{t}^{(k)}$. A long–short portfolio is then constructed by going long the top $N$ ranked stocks and short the bottom $N$ ranked stocks, resulting in a market-neutral portfolio that captures the cross-sectional signal extracted by the model.
Performance evaluation. The daily returns of the long–short portfolio are aggregated over the out-of-sample period to assess predictive performance. The primary evaluation metric is the Sharpe ratio, defined as the mean portfolio return divided by its standard deviation, which measures the risk-adjusted effectiveness of the proposed prediction and ranking pipeline.
II-B Original Dataset and Stock Universe
II-B1 Data source and time span
@@
II-B2 Universe definition and Universe0
@@
II-C Key Fields and Financial Interpretation
II-C1 Price data (stock_prices)
The main price dataset, available in both the training and supplemental sets, includes the fields summarized in TABLE II.
@@ II-C1 Price data (stock_prices)
II-C2 Static stock information (stock_list)
The stock_list table contains relatively static attributes for each stock, as summarized in Table III.
@@ II-C2 Static stock information (stock_list)
II-C3 Adjusted close price
@@
II-D Definition and Interpretation of the Target
@@
II-E Derived Features and Preprocessing
II-E1 Single-stock time-series features
@@
II-E2 Cross-sectional normalization
@@
II-F Label Construction and Weakly Supervised Training Samples
@@
II-G Temporal Split and Sequence Construction
II-G1 Temporal train–test split
@@
II-G2 Sequence construction for time-series models
@@
II-H Subset Sampling of Trading Days
@@ II-H1 Stride sampling
II-H2 Year-wise stratified fraction sampling
@@
II-H3 Relation between subsets and the full dataset
@@ III-E Quantum Temporal Convolutional Neural Network (QTCNN)
The central methodological contribution of this work is the Quantum Temporal Convolutional Neural Network (QTCNN)(shown as FIGURE 2 ), which addresses two key limitations of existing quantum approaches for sequential data:
@@
[…]
@@ III-F5 Training Objective
III-G Ablation Architectures
To isolate the contribution of individual QTCNN components, we introduce two ablation variants.
III-G1 TConv-MLP: Classical-Only Baseline
To determine whether the quantum circuit contributes beyond the temporal encoder, we construct a purely classical ablation model, TConv-MLP, which retains the identical temporal CNN encoder $f_{\text{TConv}}$ from QTCNN but replaces the quantum circuit entirely with a direct classical pathway:
$\hat{y}_{\text{TConv-MLP}}=\sigma\!\left(\text{MLP}(\bm{z}_{\text{QTCNN}})\right),\quad\bm{z}_{\text{QTCNN}}=f_{\text{TConv}}(X;\,\bm{\phi})\in\mathbb{R}^{n_{q}}.$ (23)
The MLP has the same architecture as the QTCNN classification head ($64\to 32\to 1$ with ReLU), and all training hyperparameters (optimizer, learning rate, epochs, batch size) are kept identical. The only difference is the absence of the quantum convolutional and pooling layers; consequently, the MLP input is $\bm{z}\in\mathbb{R}^{n_{q}}$ rather than $q;\,\bm{z}]\in\mathbb{R}^{n_{q}+1}$. This ablation directly tests whether the quantum circuit provides representational value beyond what the temporal encoder alone can capture.
III-G2 QTCNN-NoEnt: Entanglement Ablation
To quantify the specific contribution of quantum entanglement, we construct QTCNN-NoEnt, which is identical to the full QTCNN except that all CNOT gates are removed from the quantum convolutional layers. The modified convolutional unit becomes:
$U_{\text{conv}}^{\text{NoEnt}}(\bm{\theta})$ $=\bigl(R_{Y}(\theta_{1})\otimes R_{Z}(\theta_{2})\bigr)$ (24)
$\quad\cdot\bigl(R_{Y}(\theta_{3})\otimes R_{Z}(\theta_{4})\bigr)$
$\quad\cdot\bigl(R_{Y}(\theta_{5})\otimes R_{Z}(\theta_{6})\bigr),$
which applies only local single-qubit rotations without inter-qubit coupling. The resulting circuit operates on a product state throughout, meaning the qubits remain unentangled and each feature is processed independently. This ablation directly tests whether entanglement-induced feature interactions are responsible for QTCNN’s performance advantage.
III-H Quantum Long Short-Term Memory (QLSTM)
@@ III-H1 Sequence Embedding
III-H2 Quantum Gate Computation
@@
III-H3 LSTM Update Equations
@@
III-I Variational Quantum Neural Network (QNN)
@@ III-I3 Output Layer
III-J Quantum Kernel Support Vector Machine (QSVM)
@@ III-J3 SVM Optimization
III-K Quantum-Adaptive Self-Attention (QASA)
@@ III-K1 Transformer Encoder
III-K2 Quantum Enhancement Layer
@@ III-K3 Classification Output
III-L PQC-Enhanced Gradient Boosting (PQC+LGBM)
@@ III-L1 Stage 1: PQC Embedding
III-L2 Stage 2: LambdaRank Optimization
@@
III-M Classical Baseline Models
III-M1 Multi-Layer Perceptron (MLP)
@@
III-M2 Long Short-Term Memory (LSTM)
@@ III-M4 LightGBM Ranker
III-M5 Momentum-Volatility Baseline
@@
III-N Training Configuration
III-N1 Optimization Details
@@
III-N2 Hyperparameter Summary
TABLE IV: Training Configuration Summary
@@ III-N2 Hyperparameter Summary
III-O Evaluation Framework
III-O1 Portfolio Construction
@@ III-O3 Statistical Significance
III-P Quantum Implementation Details
III-P1 Circuit Simulation
@@
III-P2 Hardware-Efficient Ansatz
@@ III-P3 Gradient Computation
III-Q Reproducibility
III-Q1 Random Seeds
@@ III-R Conclusion
III-S Computational Considerations and Scalability
A critical challenge in variational quantum algorithm research is the exponential scaling of classical simulation. For an $n_{q}$-qubit system, state vector simulation requires $O(2^{n_{q}})$ memory and $O(L\cdot 4^{n_{q}})$ time complexity per forward pass, where $L$ denotes circuit depth. However, these costs are artifacts of classical simulation and do not reflect the complexity of execution on quantum hardware.
III-S1 Quantum Hardware Execution Complexity
On a real quantum processor, a single forward pass of the QTCNN requires a gate count $G(n_{q})$ given by:
$G(n_{q})=n_{q}+6\cdot\sum_{l=1}^{L_{\text{eff}}}\lfloor n_{q}/2^{l}\rfloor$ (58)
where the first term accounts for the $R_{Y}$ gates used in angle encoding. For the QTCNN architecture, we find $G(4)=22$, $G(8)=50$, $G(12)=72$, and $G(16)=106$. Due to the hierarchical halving structure, the circuit depth $D(n_{q})$ scales as $O(\log^{2}(n_{q}))$ when considering parallel gate execution, which is sublinear in the number of qubits. Training the model requires $2p+1$ circuit evaluations per gradient step using the parameter-shift rule, where $p$ is the number of quantum parameters.
III-S2 Qubit Count Justification
In this study, we set $n_{q}=8$ for all primary experiments. This choice is motivated by several factors:
  • Feature Variance: PCA analysis of the input features shows that the first 8 principal components capture approximately 92% of the cross-sectional variance, suggesting that 8 qubits are sufficient to represent the essential financial signals.
  • Hilbert Space Capacity: An 8-qubit system provides a 256-dimensional complex Hilbert space, which is sufficiently rich to model complex multi-feature interactions.
  • Trainability: Barren plateaus worsen exponentially with qubit count; $n_{q}=8$ maintains a favorable balance between expressibility and gradient stability.
  • Hardware Compatibility: This scale is compatible with current NISQ devices, such as IBM’s Eagle (127 qubits) and Google’s Sycamore (53 qubits).
Table V reports the empirical wall-clock time for a single forward-backward pass across all quantum architectures evaluated in this study, measured on PennyLane. Note that these times reflect classical simulation overhead on a state-vector backend.
TABLE V: Computational Cost per Training Iteration (on RTX3090 GPU and 12th Gen Intel(R) Core(TM) i9-12900KS CPU). Note: Simulation costs scale exponentially, whereas hardware execution scales polynomially.
@@ IV-B Out-of-Sample Performance Analysis
Table VI presents the out-of-sample (OOS) Sharpe ratio for all evaluated models over a 22-day testing period. Among all approaches, the quantum models demonstrate a clear advantage in terms of risk-adjusted returns.
@@
TABLE VI: Out-of-sample performance summary (22 OOS days, 5 independent runs). TConv-MLP and QTCNN-NoEnt are ablation variants of QTCNN.
Type Model Sharpe Ratio (OOS) OOS Days $p$-value vs QTCNN
@@
Ablation QTCNN-NoEnt 0.451 $\pm$ 0.044 22 0.038
Ablation TConv-MLP 0.412 $\pm$ 0.039 22 0.019
Quantum QNN 0.467 $\pm$ 0.038 22 0.048
Quantum QCNN 0.361 $\pm$ 0.298 22 0.182
Quantum QLSTM 0.333 $\pm$ 0.051 22 0.003
Classical Transformer 0.313 $\pm$ 0.036 22 0.002
Quantum QASA 0.266 $\pm$ 0.045 22 <<0.001
Quantum QSVM 0.131 $\pm$ 0.129 22 0.008
Quantum PQC+LGBM 0.122 $\pm$ 0.017 22 <<0.001
Classical LSTM 0.044 $\pm$ 0.029 22 <<0.001
$p$-values from paired bootstrap test (10,000 resamples) on daily long-short returns; see Section IV-F.
@@ IV-C Ablation Analysis
The inclusion of TConv-MLP and QTCNN-NoEnt as ablation baselines allows us to decompose QTCNN’s performance advantage into three distinct components: the temporal encoder, the quantum nonlinearity (single-qubit rotations), and entanglement (CNOT gates).
IV-C1 Contribution of the Quantum Circuit
[…]
IV-C2 Contribution of Entanglement
[…]
IV-C3 Decomposition Summary
[…]
  1. 1.
    Temporal encoder: +0.099+0.099 (Transformer $\to$ TConv-MLP, 44.0% of total gain)
  2. 2.
    Quantum nonlinearity (no entanglement): +0.039+0.039 (TConv-MLP $\to$ QTCNN-NoEnt, 17.3% of total gain)
  3. 3.
    Entanglement: +0.087+0.087 (QTCNN-NoEnt $\to$ QTCNN, 38.7% of total gain)
This decomposition reveals that the quantum circuit contributes 56% of the total improvement, with entanglement alone accounting for nearly 39%. The remaining quantum contribution stems from the nonlinear feature transformation induced by single-qubit rotations in the Hilbert space.
IV-C4 Architecture-Level Comparisons